OpenStudy (anonymous):

Please help! I'll give you a medal<3 A bag contains 2 White marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. A. What is the probability of selecting a purple marble and then a white marble? B. What is the probability of selecting two white marbles? C. Is there a greater chance of selecting two white marbles in a row or two purple marbles in a row. Show your work.

OpenStudy (anonymous):

prob of selecting purple marb=7/9 prob of selecting white marb=2/9 purple is more like ly because ther is more